OpenStudy (anonymous):

what is the diameter of a hula hoop with a 1 5/6 ft radius

OpenStudy (jdoe0001):

hint: diameter is twice the radius

OpenStudy (anonymous):

I worked it and came out with 22/6? Would that be correct?

jimthompson5910 (jim_thompson5910):

1 5/6 = 11/6 double that to get 2*(11/6) = 22/6 make sure you reduce 22/6 as much as possible

OpenStudy (anonymous):

I reduce I got 11/3

jimthompson5910 (jim_thompson5910):

so that is the diameter as an improper fraction

jimthompson5910 (jim_thompson5910):

what is the diameter as a mixed number?

OpenStudy (anonymous):

3 2/3, thank you for your help

jimthompson5910 (jim_thompson5910):

11/3 = 3 2/3, correct you're welcome